40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6875     Entry time: Tue Jun 26 22:37:43 2012
Author: yuta 
Type: Update 
Category: IOO 
Subject: energized OMC stages 

[Koji, Yuta]

We checked that PZTs between SRM and OMC (called OMC stage 1 and 2) is working.
Now we need them to be EPICS channels because they are not connected to digital world right now.

Background:
  For the IFO alignment, what we have been doing for last 2weeks is,

1. Align Y arm to Y end green and maximize green transmission
2. Use PZT2 to maximize TRY (PZT1 is not functioning well. PZT1 Y do a little, but X totally does nothing.)
3. Align BS and X arm to maximize TRX
4. Tune BS and ITMX so that reflection from both arms overlap at AS
5. Align X end green to that we can see bright(~250 uW) TEM00 at transmission

  However, we found that something (Y arm axis or Y end green?) has drifted horizontally and can't make Y green transmission and TRY high level at same time. Because PZT1 is not functioning well, it is hard to compensate beam translation.
  So, now what we have to do is to align Y arm to IR incident beam. That means, we either have to realign Y end green or forget about maximizing green transmission. I think I will leave green as it is for a while because calibration of the beatbox is going on and I want to proceed to PRC.
  Anyway, if we align IFO to the IR incident beam, we see clipping in the AS port. From the contrast measurement last night, we thought clipping comes from somewhere between BS and AS port. So, we need PZTs between BS and AS port working.

What we did:
  1. Turned on 24P 24N power supplies(Sorensen DCS33-33E) in AUX_OMC_SOUTH rack to supply power to AUX_OMC_NORTH rack. 18P 18N cables to OMC_NORTH was unplugged and used by the beatbox, so we reconnected them.

  2. Turned on KEPCO high voltage power supply to supply 150 V to the PZT driver, but it was not functioning well. So, we currently use Aligent HP 6209B instead. Its on the OMC_NORTH rack.

  3. PZT driver output to OMC stage 1 was unplugged. So, we plugged them.

  4. Opened PZT driver (LIGO-D060287), put some signal from Piezo_Drive_in(J4 in schematic), and checked beamspot at AS port is moving. The gain from Piezo_Drive_in to the output (hv_out) was ~20.

  5. We could avoid clipping by putting some offset to OMC stage 2 (or 1) in yaw. That means, the clipping comes from after OMC stage 2.

Conclusion:
  If we can control OMC stage 1 and 2, we can avoid clipping. So, we want them to be EPICS channels.

ELOG V3.1.3-